Calculate the absolute temperature of the gas at which 3.23×10−3 molmol occupies 448 mL at 740 torr .

Express the temperature in kelvins to three significant digits.

2-

Calculate the pressure, in atmospheres, if 2.65×10−3 mol occupies 415 mL at 134 ∘C.

Express the pressure in atmospheres to three significant digits.

3-

Calculate the quantity of gas, in moles, if 127 L at 60.0 ∘C has a pressure of 11.35 kPa .

Express the amount in moles to three significant digits
Answer #1

1 According to the condition given in question We know foom ideal gas equation PV=nRT T= PV I NR. T= 740 torr X 0.4481 3.93 x3) Now given V=1971 T = 273 +60 = 333K Ps 11.35 x 103T = 331. 152 0.2014 T= 1650 K Tempesatuse 2) Now given n = 2165x16 mol V = 0.415 l T = 134+273 = 407 K P = ART P 2.65 X10 X0.n= PV RT P = 11.35k pa » R = 8.31 ek pa molk n = 11. 35 × 1271 8.31 X 333 K In = 0.521) moles

.